OpenStudy (anonymous):

What is the equation of the line passing through (0, 2) and perpendicular to the line y = 4x – 3 in slope-intercept form?

OpenStudy (amorfide):

okay so since it is perpendicular it passes the this equation at a right angle so this means the gradient of the perpendicular line is the negative reciprocal of the equations gradient given to you so y=4x-3 the coefficient of x is the gradient to get the negative reciprical you do -1/m = -1/4 -1/4 is your gradient now do y-y1=m(x-x1)

OpenStudy (sgadi):

4y+x=8

OpenStudy (amorfide):

sgadi do not give out answers without working out as this does not help in the slightest

OpenStudy (amorfide):

@sgadi

OpenStudy (anonymous):

sgadi that is the wrong answer thats not even ac hoice i have

OpenStudy (apoorvk):

slope of a line in the form "y=mx+c" is 'm'. here slope = m = 4 for a line perpendicular to 'y=mx+c' its equation is: -1 y = ---x +k m where 'k' is some constant. so substitute 'm' over here, and plug in the (0,2) which this line satisfies. find 'k'. then rewrite the equation of the perpendicular line in the correct form

OpenStudy (sgadi):

@amorfide I am giving answer so that they can workout and come to the solution. :) anyway you are right.

OpenStudy (apoorvk):

@TLC do you follow the process?

OpenStudy (sgadi):

@TLC the answer is 4y+x=8 the answer may be in another form.... like y = -(1/4)x+2

OpenStudy (anonymous):

they r not perpendiular!!! i draw them@TLC